# Shallow brambles

### v3

@@ Abstract.
G. Joret is supported by the Belgian National Fund for Scientific Research (FNRS). P. Micek is supported by the National Science Center of Poland under grant UMO-2023/05/Y/ST6/00079 within the WEAVE-UNISONO program.
1. Introduction
@@
It was pointed out to us by Michał Pilipczuk that our definition of a shallow bramble is related to the cop-width of a graph, a parameter recently introduced by Toruńczyk [18]. Note that very much like brambles (or heavens) in the helicopter version of the cops and robber game of Seymour and Thomas [16], the existence of a depth-$r$ bramble of order $k$ provides a winning strategy for the robber against $k-1$ cops in the version of the game by Toruńczyk when the robber has speed $12r+$ 1.
@@ 2. Definitions and notation
We say that a subset $B\subseteq V(G)$ is connected if the induced subgraph $G[B]$ is. A minor-model of a graph $H$ in a graph $G$ is a collection =$\mathcal{M}=\{M_{v}:v\in V(H)\}$ of pairwise disjoint connected subsets of $V(G)$, one for each vertex $v$ of $H$, such that there is an edge in $G$ between $M_{v}$ and $M_{w}$ for every edge $vw\in E(H)$. If moreover $G[M_{v}]$ has radius at most $r$ for every $v\in V(H)$, we say that $\mathcal{M}$ is a depth-$r$ minor-model of $H$ in $G$.
@@
Observe that every depth-$r$ $3$-bramble is a depth-$r$ tangle, and every depth-$r$ tangle is a depth-$r$ bramble, so we have $\mathrm{bn}_{r,3}(G)\leqslant\mathrm{tn}_{r}(G)\leqslant\mathrm{bn}_{r}(G)$ for every graph $G$.
Let $r\in\mathbb{N}\cup\{\infty\}$ and let $k$ be an integer with $k\geqslant 1$. We say that a vertex set $S$ in a graph $G$ is depth-$r$ $k$-linked if for any set $X$ of fewer than $k$ vertices, there is a connected subset in $G-X$ of radius at most $r$ that contains more than half of the vertices of $S$. The depth-$r$ linkedness of $G$, denoted $\mathrm{link}_{r}(G)$, is the largest integer $k$ for which $G$ contains a depth-$r$ $k$-linked set. The linkedness of $G$ is $\mathrm{link}_{\infty}(G)$.
We say that a set $S$ of vertices is depth-$r$ well-linked if for every two subsets $A$ and $B$ of $S$ with $|A|=|B|$ and for every vertex set $Y$ of size less than $|A|$, there exists an $A-B$ path of length $\leqslant r$ in $G-Y$. The depth-$r$ well-linkedness of $G$, denoted $\mathrm{well}_{r}(G)$, is the size of the largest depth-$r$ well-linked set in $G$. The well-linkedness of $G$ is $\mathrm{well}_{\infty}(G)$.
We note that in the case $r=\infty$, by Menger’s theorem, the notion of a depth-$r$ well-linked set could be equivalently defined by requiring the existence of $|A|$ vertex-disjoint $A-B$ paths (of length at most $r$) for every two equal-sized subsets $A$ and $B$ of $S$. However, for finite $r$ this equivalence is far from correct, due to examples in [10].
3. Main results
@@ 4. The proofs
Due to a classical result of Erdős [4], there exists a graph $G$ with minimum degree at least $d$ and girth at least $48s+$ 4. Then $\mathrm{scol}_{r}(G)\geqslant\mathrm{scol}_{1}(G)\geqslant d+1$. It remains to show that $\mathrm{bn}_{s}(G)=2$.
Let $\mathcal{B}$ be a depth-$s$ bramble in $G$. Since any two bramble elements touch and each bramble element has radius at most $s$, it follows that the graph $G(\mathcal{B})$ induced by the union of the elements of $\mathcal{B}$ has diameter at most $14s+$ 1. Since this is less than half the girth of $G$, $G(\mathcal{B})$ must induce a tree. Trees have treewidth $1$, so every bramble (in particular every depth-$s$ bramble) has a hitting set of size at most $2$ [16]. On the other hand, any edge $uv$ gives rise to the depth-$s$ bramble :=$\mathcal{B}:=\{u,v\}$ that has order $2$, so $\mathrm{bn}_{s}(G)=2$. ∎
@@
For the second inequality, let $\mathcal{B}$ be a depth-$r$ bramble of $G$ and let $\pi$ be an ordering of $V(G)$. For each bramble element $B\in\mathcal{B}$, let $v_{B}$ be the minimum vertex of $B$ w.r.t. $\pi$, and let $v^{*}$ be the maximum vertex among all $v_{B}$ w.r.t. $\pi$. More formally, $v_{B}$ is the vertex such that $\pi(v_{B})=\min_{w\in B}\pi(w)$ and $v^{*}$ is such that $\pi(v^{*})=\max_{B\in\mathcal{B}}\min_{w\in B}\pi(w)$. Since $\mathcal{B}$ is a depth-$r$ bramble, each bramble element of $\mathcal{B}$ has diameter at most $2r$. Thus for every $B\in\mathcal{B}$ there is a path $P(v^{*},B)$ from $v^{*}$ to $v_{B}$ of length at most $14r+$ 1. Traversing $P(v^{*},B)$ starting from $v^{*}$, let $x_{B}$ be the first vertex such that $\pi(x_{B})\leqslant\pi(v^{*})$. (Note that $x_{B}$ indeed exists since $\pi(v_{B})\leqslant\pi(v^{*})$.) Then $X:=\bigcup_{B\in\mathcal{B}}x_{B}$ is both a hitting set of $\mathcal{B}$ and a subset of $\mathrm{SReach}[4r+1,\pi,v^{*}]$. Since we derived this for arbitrary $\pi$ and $\mathcal{B}$, it follows that $\mathrm{bn}_{r}(G)\leqslant\mathrm{scol}_{4r+1}(G)$. ∎
@@
Note that there exists at least one good model, since a single vertex defines one. Given a minor-model $\mathcal{M}$, let $V(\mathcal{M})$ denote the union of the branch sets of $\mathcal{M}$, and let $R(\mathcal{M})$ denote the union of all bramble elements of $\mathcal{B}$ that are disjoint from $V(\mathcal{M})$. Observe that, since every two bramble elements touch, the subgraph induced by $R(\mathcal{M})$ has radius at most $13r+$ 1.
@@
The proof is done if $R(\mathcal{M})=\emptyset$, because then $V(\mathcal{M})$ is a hitting set of $\mathcal{B}$ of order at most $\sum_{1=1}^{s}|V(M_{i})|\leqslant t+(5r+1)(t-1)t\leqslant(5r+1)\cdot t^{2}$. Thus we may assume that $R(\mathcal{M})$ is non-empty, and from there we will derive a contradiction.
@@
Thanks to (1) we know that each branch set $M_{i}\in\mathcal{M}$ satisfies the following key property: There is a bramble element $B\in\mathcal{B}$ that intersects $M_{i}$ and avoids all other branch sets of $\mathcal{M}$. Take a shortest path $P_{i}$ from $M_{i}$ to the non-empty set $R(\mathcal{M})$ through $M_{i}\cup B$. Let $x_{i}$ be the endpoint of $P_{i}$ in $R(\mathcal{M})$. Observe that $P_{i}$ contains at most $12r+$ 1 edges, and thus at most $22r+$ 2 vertices. Let $Q_{i}$ be $P_{i}$ minus its endpoint in $M_{i}$ (so $Q_{i}$ contains at most $2r$ edges).
Recall that the subgraph induced by $R(\mathcal{M})$ has radius at most $13r+$ 1, so there exists a vertex $x^{*}\in R(\mathcal{M})$ and for each $1\leqslant i\leqslant s$, a path on at most $13r+$ 1 edges, included in $R(\mathcal{M})$ and connecting $x^{*}$ to $x_{i}$. Let $X\subseteq R(\mathcal{M})$ be the union of the vertex sets of these $s$ paths. Note that $X$ is connected and contains at most $1+(3r+1)s$ vertices.
Now define a new branch set $M_{s+1}$, consisting of $X$ together with the paths $Q_{1},\ldots,Q_{s}$. Then the radius of $M_{s+1}$ is at most $15r+$ 1, and $|V(M_{s+1})|\leqslant 1+(5r+1)s\leqslant 1+(5r+1)(t-1)$. Moreover, by construction $M_{s+1}$ is connected to all previous branch sets. We conclude that $\mathcal{M}+M_{s+1}$ is a good model with $|R(\mathcal{M}+M_{s+1})|<|R(\mathcal{M})|$, contradicting the minimality of $\mathcal{M}$. ∎
@@
Let $S$ be a depth-$r$ $k$-linked set, where $k:=\mathrm{link}_{r}(G)$. By definition of depth-$r$ $k$-linkedness, for each set $X\subseteq V(G)$ of order $|X|<k$, there exists an $r$-ball $B_{X}$ in $G-X$ that contains more than half of the vertices of $S$. Let $\mathcal{B}:=(B_{X})_{X\subseteq V(G),|X|<k}$ be the set of these balls. The elements of $\mathcal{B}$ pairwise intersect since each of them contains more than half of $S$, and therefore $\mathcal{B}$ is a depth-$r$ bramble of $G$. Furthermore, the order of $\mathcal{B}$ is at least $k$, since any hypothetical hitting set $X$ of order less than $k$ would not hit the bramble element $B_{X}$. Thus $\mathrm{bn}_{r}(G)\geqslant k$, as desired. ∎
@@
Since $Y$, being smaller than $S$, is not a hitting set for $\mathcal{B}$, there exists a bramble element $B\in\mathcal{B}$ which is disjoint from $Y$. Because $B$ is an element of a depth-$r$ bramble, we can choose a vertex $u$ in $B$ such that (in the graph induced by $B$ and thus also in $G-Y$) every vertex of $B$ is at distance at most $r$ from $u$. We let $U$ denote the $(3r+1)$-ball in $G-Y$ centered at $u$. It suffices to show that $U$ contains more than half of $S$, which we will now do.
Suppose for a contradiction that $|U\cap S|\leqslant\lfloor\frac{|S|}{2}\rfloor$. If there exists any, let $C\in\mathcal{B}$ be a bramble element that is disjoint from $Y$. As $S$ hits $C$, we may choose a vertex $c\in C\cap S$. Since $C$ and $B$ touch and $C$ has radius at most $r$, there is a path in $G-Y$ joining $u$ and $c$ of length at most $13r+$ 1, so that $c\in U\cap S$. In other words, $U\cap S$ hits $C$.
@@
We prove that any minimum hitting set $S$ for a depth-$r$ bramble $\mathcal{B}$ is depth-$(4r+1)$ well-linked. Suppose not, then there exist two equal-sized subsets $A,B$ of $S$ and a set $Y$ of fewer than $|A|$ vertices, such that $G-Y$ does not contain any $A-B$ path of length $\leqslant 4r+1$. Consider the sub-bramble $\mathcal{B}_{Y}$ of $\mathcal{B}$ that contains all bramble elements of $\mathcal{B}$ that are disjoint from $Y$. If both $A$ and $B$ intersect some (not necessarily the same) bramble element of $\mathcal{B}_{Y}$ then these bramble elements touch, so then there must be a path of length at most $14r+$ 1 between $A$ and $B$ in $G-Y$, a contradiction. So without loss of generality $\bigcup_{C\in\mathcal{B}_{Y}}C\cap A$ is empty. But then $(S-A)+Y$ is a hitting set for $\mathcal{B}$, contradicting the minimality of $S$. ∎
@@
Assume for a contradiction that there exists a set $X$ of fewer than $k_{S}$ vertices such that every $3r$-ball in $G-X$ contains at most half of the vertices of $S$.
We first consider the case that $G-X$ contains a $2r$-ball $B_{2r}(v)$, centered at some vertex $v$, containing more than $k_{S}$ vertices of $S$. Let $B_{3r}(v)$ be the $3r$-ball in $G-X$ centered at the same vertex. Since $|B_{3r}(v)\cap S|\leqslant\lfloor|S|/2\rfloor$, we know that $S-(X\cup B_{3r}(v))$ has size at least $|S|-(k_{S}-1)-\lfloor|S|/2\rfloor\geqslant k_{S}$. Thus we can choose a subset $A$ of $B_{2r}(v)\cap S$ and a subset $B$ of $S-(X\cup B_{3r}(v))$ such that $|A|=|B|=k_{S}>|X|$. However, by construction there cannot exist any $A-B$ path of length at most $r$ in $G-X$, contradicting our assumption that $S$ is depth-$r$ well-linked.
We have thus reduced to the case that no $2r$-ball in $G-X$ contains more than $k_{S}$ vertices of $S$.
We now greedily build a collection of $r$-balls in $G-X$, collecting at least one new vertex from $S$ at each step, and stopping as soon as the corresponding $2r$-balls (again in $G-X$ and with the same centers) cover at least half of $S$. Let $U_{r}$ denote the union of the $r$-balls, and let $U_{2r}$ denote the union of the corresponding $2r$-balls.
@@
Thus we can choose a subset $A$ of $U_{r}\cap S$ and a subset $B$ of $S-(X\cup U_{2r})$ such that $|A|=|B|=k_{S}>|X|$. However, as in the first case, every $A-B$ path in $G-X$ must have length more than $r$, contradicting that $S$ is depth-$r$ well-linked. ∎
@@
The lower bound on $\mathrm{bn}_{r}(G)$ is true because every depth-$r$ $t$-bramble is a depth-$r$ bramble, so we only need to prove the upper bound. Let $\mathcal{B}$ be a depth-$r$ bramble of order $k:=\mathrm{bn}_{r}(G)$. For each $X\subseteq V(G)$ with $|X|<\lceil\tfrac{k}{t}\rceil$, we can choose a bramble element $B_{X}$ of $\mathcal{B}$ that is disjoint from $X$, since $X$ is too small to be a hitting set of $\mathcal{B}$. Let $v_{X}$ be a vertex of $B_{X}$ such that $B_{X}$ is contained in the $r$-ball centered at $v_{X}$. Now define $T_{X}$ as the $(3r+1)$-ball in $G-X$ centered at $v_{X}$. We claim that $\mathcal{T}:=\left(T_{X}\right)_{X\subseteq V(G),|X|<\tfrac{k}{t}}$ is a depth-$(3r+1)$ $t$-bramble of $G$ of order at least $\lceil\tfrac{k}{t}\rceil$. (This then implies that $k\leqslant t\cdot\mathrm{bn}_{3r+1,t}(G)$, as desired.)
Indeed, consider $t$ sets $X_{1},\ldots,X_{t}$ of size smaller than $\lceil\tfrac{k}{t}\rceil$. Since $|\bigcup_{1\leqslant i\leqslant t}X_{i}|<k$, there is an element $B$ of $\mathcal{B}$ that is disjoint from $\bigcup_{1\leqslant i\leqslant t}X_{i}$. For each $1\leqslant i\leqslant t$, because $B_{X_{i}}$ and $B$ touch, are disjoint from $X_{i}$ and both have radius at most $r$, we see that every vertex of $B$ can be reached from $v_{X_{i}}$ by some path in $G-X_{i}$ of length at most $13r+$ 1, so that $B\subseteq T_{X_{i}}$. Thus $B\subseteq\bigcap_{1\leqslant i\leqslant t}T(X_{i})$. In particular, the intersection of $T(X_{1}),\ldots,T(X_{t})$ is non-empty, so $\mathcal{T}$ is indeed a depth-$(3r+1)$ $t$-bramble. Finally, $\mathcal{T}$ does not have any hitting set $Y$ of size strictly smaller than $\lceil\tfrac{k}{t}\rceil$, because $Y$ would not hit the element $T_{Y}$. So $\mathcal{T}$ has order at least $\lceil\tfrac{k}{t}\rceil$. ∎
5. Final remarks

### v4

@@ Abstract
keywords:
Sparsity, shallow minors, brambles, polynomial expansion
1 Introduction
@@
It was pointed out to us by Michał Pilipczuk that our definition of a shallow bramble is related to the cop-width of a graph, a parameter recently introduced by Toruńczyk [18]. Note that very much like brambles (or heavens) in the helicopter version of the cops and robber game of Seymour and Thomas [16], the existence of a depth-$r$ bramble of order $k$ provides a winning strategy for the robber against k1k-1 cops in the version of the game by Toruńczyk when the robber has speed $2r+1$.
@@ 2 Definitions and notation
We say that a subset $B\subseteq V(G)$ is connected if the induced subgraph $G[B]$ is. A minor-model of a graph $H$ in a graph $G$ is a collection $\mathcal{M}=\{M_{v}:v\in V(H)\}$ of pairwise disjoint connected subsets of $V(G)$, one for each vertex $v$ of $H$, such that there is an edge in $G$ between $M_{v}$ and $M_{w}$ for every edge $vw\in E(H)$. If moreover $G[M_{v}]$ has radius at most $r$ for every $v\in V(H)$, we say that $\mathcal{M}$ is a depth-$r$ minor-model of $H$ in $G$.
@@
Observe that every depth-$r$ 33-bramble is a depth-$r$ tangle, and every depth-$r$ tangle is a depth-$r$ bramble, so we have $\mathrm{bn}_{r,3}(G)\leqslant\mathrm{tn}_{r}(G)\leqslant\mathrm{bn}_{r}(G)$ for every graph $G$.
Let $r\in\mathbb{N}\cup\{\infty\}$ and let $k$ be an integer with $k\geqslant 1$. We say that a vertex set $S$ in a graph $G$ is depth-$r$ $k$-linked if for any set $X$ of fewer than $k$ vertices, there is a connected subset in GXG-X of radius at most $r$ that contains more than half of the vertices of $S$. The depth-$r$ linkedness of $G$, denoted $\mathrm{link}_{r}(G)$, is the largest integer $k$ for which $G$ contains a depth-$r$ $k$-linked set. The linkedness of $G$ is $\mathrm{link}_{\infty}(G)$.
We say that a set $S$ of vertices is depth-$r$ well-linked if for every two subsets $A$ and $B$ of $S$ with $|A|=|B|$ and for every vertex set $Y$ of size less than $|A|$, there exists an ABA-B path of length $\leqslant r$ in GYG-Y. The depth-$r$ well-linkedness of $G$, denoted $\mathrm{well}_{r}(G)$, is the size of the largest depth-$r$ well-linked set in $G$. The well-linkedness of $G$ is $\mathrm{well}_{\infty}(G)$.
We note that in the case $r=\infty$, by Menger’s theorem, the notion of a depth-$r$ well-linked set could be equivalently defined by requiring the existence of $|A|$ vertex-disjoint ABA-B paths (of length at most $r$) for every two equal-sized subsets $A$ and $B$ of $S$. However, for finite $r$ this equivalence is far from correct, due to examples in [10].
3 Main results
@@ 4 The proofs
Due to a classical result of Erdős [4], there exists a graph $G$ with minimum degree at least $d$ and girth at least $8s+4$. Then $\mathrm{scol}_{r}(G)\geqslant\mathrm{scol}_{1}(G)\geqslant d+1$. It remains to show that $\mathrm{bn}_{s}(G)=2$.
Let $\mathcal{B}$ be a depth-$s$ bramble in $G$. Since any two bramble elements touch and each bramble element has radius at most $s$, it follows that the graph $G(\mathcal{B})$ induced by the union of the elements of $\mathcal{B}$ has diameter at most $4s+1$. Since this is less than half the girth of $G$, $G(\mathcal{B})$ must induce a tree. Trees have treewidth $1$, so every bramble (in particular every depth-$s$ bramble) has a hitting set of size at most $2$ [16]. On the other hand, any edge $uv$ gives rise to the depth-$s$ bramble $\mathcal{B}:=\{u,v\}$ that has order $2$, so $\mathrm{bn}_{s}(G)=2$. ∎
@@
For the second inequality, let $\mathcal{B}$ be a depth-$r$ bramble of $G$ and let $\pi$ be an ordering of $V(G)$. For each bramble element $B\in\mathcal{B}$, let $v_{B}$ be the minimum vertex of $B$ w.r.t. $\pi$, and let $v^{*}$ be the maximum vertex among all $v_{B}$ w.r.t. $\pi$. More formally, $v_{B}$ is the vertex such that $\pi(v_{B})=\min_{w\in B}\pi(w)$ and $v^{*}$ is such that $\pi(v^{*})=\max_{B\in\mathcal{B}}\min_{w\in B}\pi(w)$. Since $\mathcal{B}$ is a depth-$r$ bramble, each bramble element of $\mathcal{B}$ has diameter at most $2r$. Thus for every $B\in\mathcal{B}$ there is a path $P(v^{*},B)$ from $v^{*}$ to $v_{B}$ of length at most $4r+1$. Traversing $P(v^{*},B)$ starting from $v^{*}$, let $x_{B}$ be the first vertex such that $\pi(x_{B})\leqslant\pi(v^{*})$. (Note that $x_{B}$ indeed exists since $\pi(v_{B})\leqslant\pi(v^{*})$.) Then $X:=\bigcup_{B\in\mathcal{B}}x_{B}$ is both a hitting set of $\mathcal{B}$ and a subset of $\mathrm{SReach}[4r+1,\pi,v^{*}]$. Since we derived this for arbitrary $\pi$ and $\mathcal{B}$, it follows that $\mathrm{bn}_{r}(G)\leqslant\mathrm{scol}_{4r+1}(G)$. ∎
@@
Note that there exists at least one good model, since a single vertex defines one. Given a minor-model $\mathcal{M}$, let $V(\mathcal{M})$ denote the union of the branch sets of $\mathcal{M}$, and let $R(\mathcal{M})$ denote the union of all bramble elements of $\mathcal{B}$ that are disjoint from $V(\mathcal{M})$. Observe that, since every two bramble elements touch, the subgraph induced by $R(\mathcal{M})$ has radius at most $3r+1$.
@@
The proof is done if $R(\mathcal{M})=\emptyset$, because then $V(\mathcal{M})$ is a hitting set of $\mathcal{B}$ of order at most
$\sum_{1=1}^{s}|V(M_{i})|\leqslant t+(5r+1)(t-1)t\leqslant(5r+1)\cdot t^{2}.$
Thus we may assume that $R(\mathcal{M})$ is non-empty, and from there we will derive a contradiction.
@@
Thanks to (1) we know that each branch set $M_{i}\in\mathcal{M}$ satisfies the following key property: There is a bramble element $B\in\mathcal{B}$ that intersects $M_{i}$ and avoids all other branch sets of $\mathcal{M}$. Take a shortest path $P_{i}$ from $M_{i}$ to the non-empty set $R(\mathcal{M})$ through $M_{i}\cup B$. Let $x_{i}$ be the endpoint of $P_{i}$ in $R(\mathcal{M})$. Observe that $P_{i}$ contains at most $2r+1$ edges, and thus at most $2r+2$ vertices. Let $Q_{i}$ be $P_{i}$ minus its endpoint in $M_{i}$ (so $Q_{i}$ contains at most $2r$ edges).
Recall that the subgraph induced by $R(\mathcal{M})$ has radius at most $3r+1$, so there exists a vertex $x^{*}\in R(\mathcal{M})$ and for each $1\leqslant i\leqslant s$, a path on at most $3r+1$ edges, included in $R(\mathcal{M})$ and connecting $x^{*}$ to $x_{i}$. Let $X\subseteq R(\mathcal{M})$ be the union of the vertex sets of these $s$ paths. Note that $X$ is connected and contains at most $1+(3r+1)s$ vertices.
Now define a new branch set $M_{s+1}$, consisting of $X$ together with the paths $Q_{1},\ldots,Q_{s}$. Then the radius of $M_{s+1}$ is at most $5r+1$, and $|V(M_{s+1})|\leqslant 1+(5r+1)s\leqslant 1+(5r+1)(t-1)$. Moreover, by construction $M_{s+1}$ is connected to all previous branch sets. We conclude that $\mathcal{M}+M_{s+1}$ is a good model with $|R(\mathcal{M}+M_{s+1})|<|R(\mathcal{M})|$, contradicting the minimality of $\mathcal{M}$. ∎
@@
Let $S$ be a depth-$r$ $k$-linked set, where $k:=\mathrm{link}_{r}(G)$. By definition of depth-$r$ $k$-linkedness, for each set $X\subseteq V(G)$ of order $|X|<k$, there exists an $r$-ball $B_{X}$ in GXG-X that contains more than half of the vertices of $S$. Let $\mathcal{B}:=(B_{X})_{X\subseteq V(G),|X|<k}$ be the set of these balls. The elements of $\mathcal{B}$ pairwise intersect since each of them contains more than half of $S$, and therefore $\mathcal{B}$ is a depth-$r$ bramble of $G$. Furthermore, the order of $\mathcal{B}$ is at least $k$, since any hypothetical hitting set $X$ of order less than $k$ would not hit the bramble element $B_{X}$. Thus $\mathrm{bn}_{r}(G)\geqslant k$, as desired. ∎
@@
Since $Y$, being smaller than $S$, is not a hitting set for $\mathcal{B}$, there exists a bramble element $B\in\mathcal{B}$ which is disjoint from $Y$. Because $B$ is an element of a depth-$r$ bramble, we can choose a vertex $u$ in $B$ such that (in the graph induced by $B$ and thus also in GYG-Y) every vertex of $B$ is at distance at most $r$ from $u$. We let $U$ denote the $(3r+1)$-ball in GYG-Y centered at $u$. It suffices to show that $U$ contains more than half of $S$, which we will now do.
Suppose for a contradiction that $|U\cap S|\leqslant\lfloor\frac{|S|}{2}\rfloor$. If there exists any, let $C\in\mathcal{B}$ be a bramble element that is disjoint from $Y$. As $S$ hits $C$, we may choose a vertex $c\in C\cap S$. Since $C$ and $B$ touch and $C$ has radius at most $r$, there is a path in GYG-Y joining $u$ and $c$ of length at most $3r+1$, so that $c\in U\cap S$. In other words, $U\cap S$ hits $C$.
@@
We prove that any minimum hitting set $S$ for a depth-$r$ bramble $\mathcal{B}$ is depth-$(4r+1)$ well-linked. Suppose not, then there exist two equal-sized subsets $A,B$ of $S$ and a set $Y$ of fewer than $|A|$ vertices, such that GYG-Y does not contain any ABA-B path of length $\leqslant 4r+1$. Consider the sub-bramble $\mathcal{B}_{Y}$ of $\mathcal{B}$ that contains all bramble elements of $\mathcal{B}$ that are disjoint from $Y$. If both $A$ and $B$ intersect some (not necessarily the same) bramble element of $\mathcal{B}_{Y}$ then these bramble elements touch, so then there must be a path of length at most $4r+1$ between $A$ and $B$ in GYG-Y, a contradiction. So without loss of generality $\bigcup_{C\in\mathcal{B}_{Y}}C\cap A$ is empty. But then (SA)+Y(S-A)+Y is a hitting set for $\mathcal{B}$, contradicting the minimality of $S$. ∎
@@
Assume for a contradiction that there exists a set $X$ of fewer than $k_{S}$ vertices such that every $3r$-ball in GXG-X contains at most half of the vertices of $S$.
We first consider the case that GXG-X contains a $2r$-ball $B_{2r}(v)$, centered at some vertex $v$, containing more than $k_{S}$ vertices of $S$. Let $B_{3r}(v)$ be the $3r$-ball in GXG-X centered at the same vertex. Since $|B_{3r}(v)\cap S|\leqslant\lfloor|S|/2\rfloor$, we know that $S-(X\cup B_{3r}(v))$ has size at least $|S|-(k_{S}-1)-\lfloor|S|/2\rfloor\geqslant k_{S}$. Thus we can choose a subset $A$ of $B_{2r}(v)\cap S$ and a subset $B$ of $S-(X\cup B_{3r}(v))$ such that $|A|=|B|=k_{S}>|X|$. However, by construction there cannot exist any ABA-B path of length at most $r$ in GXG-X, contradicting our assumption that $S$ is depth-$r$ well-linked.
We have thus reduced to the case that no $2r$-ball in GXG-X contains more than $k_{S}$ vertices of $S$.
We now greedily build a collection of $r$-balls in GXG-X, collecting at least one new vertex from $S$ at each step, and stopping as soon as the corresponding $2r$-balls (again in GXG-X and with the same centers) cover at least half of $S$. Let $U_{r}$ denote the union of the $r$-balls, and let $U_{2r}$ denote the union of the corresponding $2r$-balls.
@@
Thus we can choose a subset $A$ of $U_{r}\cap S$ and a subset $B$ of $S-(X\cup U_{2r})$ such that $|A|=|B|=k_{S}>|X|$. However, as in the first case, every ABA-B path in GXG-X must have length more than $r$, contradicting that $S$ is depth-$r$ well-linked. ∎
@@
The lower bound on $\mathrm{bn}_{r}(G)$ is true because every depth-$r$ $t$-bramble is a depth-$r$ bramble, so we only need to prove the upper bound. Let $\mathcal{B}$ be a depth-$r$ bramble of order $k:=\mathrm{bn}_{r}(G)$. For each $X\subseteq V(G)$ with $|X|<\lceil\tfrac{k}{t}\rceil$, we can choose a bramble element $B_{X}$ of $\mathcal{B}$ that is disjoint from $X$, since $X$ is too small to be a hitting set of $\mathcal{B}$. Let $v_{X}$ be a vertex of $B_{X}$ such that $B_{X}$ is contained in the $r$-ball centered at $v_{X}$. Now define $T_{X}$ as the $(3r+1)$-ball in GXG-X centered at $v_{X}$. We claim that $\mathcal{T}:=\left(T_{X}\right)_{X\subseteq V(G),|X|<\tfrac{k}{t}}$ is a depth-$(3r+1)$ $t$-bramble of $G$ of order at least $\lceil\tfrac{k}{t}\rceil$. (This then implies that $k\leqslant t\cdot\mathrm{bn}_{3r+1,t}(G)$, as desired.)
Indeed, consider $t$ sets $X_{1},\ldots,X_{t}$ of size smaller than $\lceil\tfrac{k}{t}\rceil$. Since $|\bigcup_{1\leqslant i\leqslant t}X_{i}|<k$, there is an element $B$ of $\mathcal{B}$ that is disjoint from $\bigcup_{1\leqslant i\leqslant t}X_{i}$. For each $1\leqslant i\leqslant t$, because $B_{X_{i}}$ and $B$ touch, are disjoint from $X_{i}$ and both have radius at most $r$, we see that every vertex of $B$ can be reached from $v_{X_{i}}$ by some path in $G-X_{i}$ of length at most $3r+1$, so that $B\subseteq T_{X_{i}}$. Thus $B\subseteq\bigcap_{1\leqslant i\leqslant t}T(X_{i})$. In particular, the intersection of $T(X_{1}),\ldots,T(X_{t})$ is non-empty, so $\mathcal{T}$ is indeed a depth-$(3r+1)$ $t$-bramble. Finally, $\mathcal{T}$ does not have any hitting set $Y$ of size strictly smaller than $\lceil\tfrac{k}{t}\rceil$, because $Y$ would not hit the element $T_{Y}$. So $\mathcal{T}$ has order at least $\lceil\tfrac{k}{t}\rceil$. ∎
5 Final remarks